# Open-World Evaluation for Retrieving Diverse Perspectives

### v1

@@ Appendix A Identifying whether a perspective is represented in a document
Our primary goal is developing and evaluating retrieval methods for uncovering diverse perspectives. To this end, we need a system that can identify whether the retrieved document set contains a perspective. Unlike simple factoid QA setting, where we can evaluate whether a document contains an answer by string match Lee et al. (2019); Min et al. (2021a, b), evaluating whether a document contains a perspective is nontrivial, a research question on its own Sen et al. (2018); AlDayel and Magdy (2021). In this section, we aim to build an efficient model for this subtask, which is a necessary component for the evaluation of diverse retrieval models.
@@
For each document-perspective pair, we consider human judgment as the gold label.
The human-labeled test set contains 542 document-perspective pairs, which are annotated by the authors of this paper. We compute inter-annotator agreement on 176 examples, yielding a Cohen Kappa agreement of 0.56. 16.8% of the data receive a positive label. The examples are taken from top five retrieval results from the Sphere corpus with BM25 Robertson et al. (2009) and DPR Karpukhin et al. (2020a) retriever described in Section 5, and equally split between three datasets. Details on the construction of the human reference set can be found in Section B.4.
A.1 Evaluated Approaches
@@
Instead of using an NLI model, we can prompt large-scale LLM to decide whether a document contains a perspective. From the open-source LMs, we choose Llama-2-chat Touvron et al. (2023), an instruct fine-tuned version of Mistral-7B Jiang et al. (2023), Zephyr-7B Tunstall et al. (2023), and Gemma-7B Team et al. (2024). Lastly, we use GPT-4 (gpt-4-0613). To reach a better absolute performance, we additionally experiment with prompting with one in-context example for GPT-4.
@@
Prompting GPT4 showed promising results in this task, but it would require additional API calls, which incur extra cost, whenever we would like to evaluate new systems on BeRDS. We thus fine-tune the Mistral-7B model with GPT-4 predictions with one in-context example as labels. The fine-tuning dataset contains 3K examples corresponding to 50 unique questions, and implementation details can be found in Section B.5.
A.2 Results
@@
As the fine-tuned Mistral-7B comes close to GPT-4 performance, we use this model as the automatic metric for evaluating retrieval diversity, providing an open-source model that enables reproducible, cheaper evaluation.
Appendix B Experimental Details
@@
B.1 Prompts and Examples for Data Generation
Prompts for generating each dataset are shown in Tables 9, 10 and 12. Examples showing the generated data and data taken from the source can be found in Table 11.
B.2 Prompts for Perspective Detection
@@
B.3 Prompts for Query-Expansion
@@
B.4 Details on Constructing Perspective Detection Test Set
@@
B.5 Details of Fine-tuning Mistral-7B Models
We fine-tune the Mistral-7B model with LoRA Hu et al. (2021) on eight NVIDIA A40 GPUs. We train the model on 3K examples of GPT-4 labeled data on perspective detection. The training dataset is constructed using the same process as described in Section B.4, except on 50 questions, yielding five times more data.
@@ B.5 Details of Fine-tuning Mistral-7B Models
B.6 Detail of Re-ranking Experiments
@@
B.7 Examples of Generated Perspectives for Query Expansion
@@
B.8 Details of Determining Supporting and Opposing Perspectives
@@ Appendix C Additional Results
Google Search BM25 33.0 36.40 34.92 38.50 33.81 29.25 33.91 34.72
Output Contriever 48.65 44.53 47.95 44.32 42.73 35.49 46.44 41.45
@@
C.1 Results for Top 10 Documents

### v2

@@ 5 Perspective Detection
We aim to build evalution metrics for retrieval diversity. Unlike factoid QA, where we can evaluate whether a document contains an answer by string match Lee et al. (2019); Min et al. (2021a, b), evaluating whether a document contains a perspective is nontrivial, a research question on its own Sen et al. (2018); AlDayel and Magdy (2021). We need a system that identifies whether the retrieved document set contains a perspective. We define the subtask as “Perspective Detection”, and detail the process of building an efficient model for this subtask.
@@
For each document-perspective pair, we consider human judgment as the gold label. The human-labeled test set contains 542 document-perspective pairs, which are annotated by the authors of this paper. We compute inter-annotator agreement on 176 examples, yielding a Cohen Kappa agreement of 0.56. 16.8% of the data receive a positive label. The examples are taken from top five retrieval results from the Sphere corpus with BM25 Robertson et al. (2009) and DPR Karpukhin et al. (2020a) retriever described in Section 6, and equally split between three datasets. Details on the construction of the human reference set can be found in Section A.4.
5.1 Evaluated Approaches
@@
Instead of using an NLI model, we can prompt LLMs to decide whether a document contains a perspective. From the open-source LMs, we choose Llama-2-chat Touvron et al. (2023), an instruct-tuned version of Mistral-7B Jiang et al. (2023), Zephyr-7B Tunstall et al. (2023), and Gemma-7B Team et al. (2024). Lastly, we use GPT-4 (gpt-4-0613). For GPT-4, the best-performing model, we further experiment with prompting with one in-context example.
@@
Prompting GPT-4 shows promising results in this task but requires additional API calls, which incur extra costs whenever we want to evaluate new systems on BeRDS. We thus fine-tune the Mistral-7B model with GPT-4 predictions with one in-context example as labels. The fine-tuning dataset contains 3K examples corresponding to 50 unique questions, and implementation details can be found in Section A.5.
5.2 Results
@@
As the fine-tuned Mistral-7B comes close to GPT-4 performance, we use this model as the automatic metric for evaluating retrieval diversity, providing reproducible, cheaper evaluation.
6 Retrieval Approaches
We consider a suite of retrievers (sparse and dense) on our dataset. As these retrievers do not encourage the diversity of their output, we implement a re-ranking method and a query expansion method.
Base Models
We consider one sparse retriever (BM25) and four dense retrievers, and we describe them below.
  • BM25 Robertson et al. (2009): a widely adopted bag-of-word retrieval function considering token-matching between questions and documents. We use the Pyserini package for the Wikipedia corpus. For the Sphere corpus, we use the index provided by  Piktus et al. (2021).
  • DPR Karpukhin et al. (2020a): a dual encoder trained with contrastive learning on the Natural Questions (NQ) dataset Kwiatkowski et al. (2019).
  • Contriever Izacard et al. (2022): an encoder trained with a self-supervised contrastive objective. [^8] DPR uses separate encoders for document and question embeddings, and Contriever uses a shared encoder. Both models are initialized from a BERT-base encoder. We use the version that is further fine-tuned on the MSMARCO dataset Nguyen et al. (2016) after self-supervised learning, which performed best in the original paper.
  • Tart Asai et al. (2023): an encoder trained to make use of instruction in addition to the query. We implement Tart-full, the cross-encoder version, and the top hundred documents from Contriever are used as initial candidates.
  • NV-Embed-v2 Lee et al. (2024): the state-of-the-art encoder. [^9] Ranked first in the MTEB Muennighoff et al. (2022) benchmark as of Oct 15, 2024. We do not include results on Sphere due to constraints on the compute. Other BERT-based encoder models (110M parameters) uses an embedding size of 768, while NV-Embed-v2 model (7B parameters) uses an embedding size of 4096, significantly increasing the cost of building document indices.
Re-ranking
where $R$ denotes a set of top retrieved documents considered for re-ranking ($|R|=100$ in our case), $S$ denotes the document set that is already selected, and $\lambda$ is a hyperparameter, which we tune on the development split. [^10] We normalize the $Sim_{1}$ scores to be between 0 and 1, and consider $\lambda$ values of [0.5, 0.75, 0.9, 0.95, 0.99]. $Sim_{1}$ is the score from the retriever. $Sim_{2}$ is the cosine similarity between two document embeddings computed using the Universal AnglE Embedding Li and Li (2023). We choose this model as it is the top-ranked open model on the MTEB leaderboard Muennighoff et al. (2023) with less than 1B parameters (335M). Compact models are preferable as we compute document embeddings for the top hundred documents per example. The experimental details can be found in Section A.6.
Query Expansion
[…]
Corpus Retriever Default + Re-ranking + Query Expansion
Prec. MRec. Prec. MRec. Prec. MRec.
DPR 17.33 6.21 16.62 (-4.1%) 6.06 (-2.5%) 17.78 (+2.6%) 6.65 (+7.1%)
Table 5: Performances on BeRDS after re-ranking and query expansion. Each cell reports the macro-average over three datasets in BeRDS. Bolded numbers are the best performance in each row. Blue numbers indicate increases compared to the default setting.
7 Experimental Results
We report MRecall @ $5$ and Precision @ $5$, [^12] We additionally report the results for $k$ = 10 in Section B.1, where very similar trends hold. as defined in Section 3. All evaluation metrics require judging whether each retrieved document contains a perspective, or perspective detection (Section 5). We use a fine-tuned Mistral-7B model as the evaluator. [^13] We will open-source this model to promote future work.
Table 3 presents the performance of base retrievers on BeRDS using three retrieval corpus described in Section 4.2. We do not report DPR results on Google Search outputs as we find it does not generalize well beyond the Wikipedia corpus. We do not report NV-Embed-v2 results on Sphere due to limited compute.
Effect of Corpus.
Comparing retrievers operating on Wikipedia and Sphere, most retrievers achieve higher MRecall and Precision when retrieving from Sphere, except DPR. We hypothesize that DPR is trained only on the documents from the Wikipedia corpus so that it is not generalizing to the web corpus. Taking web search results (Google Search) as a corpus shows improved diversity over Wikipedia, but lags behind retrieving from the entire web (Sphere). The gap in the corpus size might cause this – we only take the top 100 Google Search results and build corpora with them (resulting in a corpus size of 4.4K passages on average), while Sphere contains 906M passages.
Comparing Retrievers.
Across all experimental settings, Tart achieves the highest diversity, followed by Contriever, BM25, and DPR. BM25 generalizes better than DPR, which is fine-tuned on NQ with Wikipedia corpus. Tart achieves higher diversity than Contriever, suggesting that specifying the task with instructions helps diversity. Even the best setting (Tart on Sphere) only achieves an average MRecall of 32.8. This indicates even the best retriever struggles to retrieve a comprehensive set of documents. We investigate the number of documents needed to be retrieved to cover diverse perspectives in Section B.2. Precision @ 1 is as high as 58.67, showing that retrievers could surface one of the perspectives.
Results: Approaches to Improve Diversity
We report the results of re-ranking and query expansion approaches in Table 5. We report the results with three retrievers (BM25, DPR, Contriever), excluding TART and NV-Embed2 for computational costs. Re-ranking improves diversity on four of eight retriever-corpus settings, at the cost of Precision for all settings. The limited performance gains could be because semantic dissimilarity does not always infer differing perspectives. Query expansion improves the Precision for five settings and boosts MRecall for six settings (over 10% for three of them). This indicates that GPT-4 could diversify the search space by providing queries hinting different perspectives to the retrievers. Future work can explore combining both approaches.
8 Analysis
Retrievers often fail to surface documents that cover comprehensive perspectives. Can we characterize the perspective that is more likely to be surfaced? We focus on questions with only two perspectives, one taking the supporting stance towards the question and the other opposing. [^14] We identify this by prompting GPT-4; details and prompts can be found in Section A.8. We report whether the top five documents obtained by the retrievers cover only supporting, opposing, both, or neither perspectives in Figure 2. We observe that all retrievers retrieve the supporting perspectives significantly more often when they fail to retrieve both.
BM.DP.Co.Ta.BM.DP.Co.Ta.BM.Co.Ta.000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Wiki     Sphere    Google APIbothnonefirstsecond
Figure 2: We report whether retrievers prefer the supporting or the opposing perspectives. Results are computed on the top-5 documents aggregated over three datasets. All retrievers favor the supporting perspectives. BM. denotes BM25 retriever, DP. denotes DPR, and Co. denotes Contriever. Ta. denotes Tart and Wiki. denotes the Wikipedia corpus.
Retriever Sycophancy.
We quantify leaning towards supporting perspectives to be $\Delta$ = ($p$ $n$)/$p$, where $p$ and $n$ are the percentages of documents containing supporting and opposing perspectives, respectively. Greater $\Delta$ indicates the retrievers favor the supporting perspectives more. We present the results on the top 5 documents retrieved from the Wikipedia corpus in Figure 3.151515Results on the other corpora show the same trends, except for DPR on the web corpus. Querying the retrievers with supporting perspectives increases $\Delta$ across the board, and retrieving with opposing perspectives decreases $\Delta$. Retrievers tend to favor perspectives that they are prompted with. This partially explains the gains from the query expansion.
0010101010202020203030303040404040505050506060606070707070Wiki_BM25Wiki_DPRWiki_Cont.Wiki_Tartsupporting perspectivedefault queryopposing perspective
Is retrieval diversity limited by the corpora?
Unlike prior work, we do not assume a gold corpus that contains target information. The best MRecall @ $5$ on BeRDS achieved by any approach on average is below 40. How much of the failure is caused by the limitation of the corpus or that of retrievers? We answer this question by computing MRecall over a larger document set, where we combine the top 100 retrieved documents from all retrievers except NV-Embed-v2 for each corpus. If this number nears 100, there exists an optimal subset of $k$ documents for each question that achieve MRecall @ $k$ close to 1. We present the results in Table 6. Only the Wikipedia corpus contains insufficiently diverse information, achieving MRecall of 62% when combining retriever outputs. In contrast, the web corpora do not limit retrievers’ performances in diversity. This result hints at great room for improvement for the retrievers.
Corpus Arguana Kialo OpinionQA Average
Wiki. 71.87 61.24 55.10 62.74
Sphere 92.80 85.27 89.46 89.18
Google 82.67 82.69 77.66 81.01
Table 6: MRecall over the union of the top 100 retrieved documents from all retrievers except NV-Embed-v2 for each corpus. These values are substantially higher than MRecall @ $5$, suggesting the corpora (especially Sphere) have good information coverage.
9 Conclusion
We study retrieval diversity for complex, subjective questions. We curate a benchmark BeRDS that evaluates retrieval diversity, provides automatic metrics, and establishes the performances of retriever baselines. Our experiments show that current retrievers cannot retrieve comprehensive information in the proposed setting.
After retrieving diverse documents, summarizing such rich information into a coherent response Laban et al. (2022); Huang et al. (2023); Zhang et al. (2023b) could be useful for the users.
Limitations
  • The perspective set in our dataset is mostly binary (the average number of perspectives is 2.9 for Kialo and 2.0 for Arguana and OpinionQA). More fine-grained perspectives could be explored (e.g. reasons for supporting positive stance).
Acknowledgements
We thank the members of UT NLP, especially Fangyuan Xu, for their feedback on the draft. The project is partially supported by a grant from Open Philanthropy and NSF grant RI-2312948. We also thank Chitrank Gupta for contributing to the earlier phase of this work. We also thank Ayon Das, an undergraduate student at UT Austin, to help with perspective detection annotation. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.
Appendix A Experimental Details
@@
A.1 Prompts and Examples for Data Generation
Prompts for generating each dataset are shown in Tables 8, 9 and 11. Examples showing the generated data and data taken from the source can be found in Table 10.
A.2 Prompts for Perspective Detection
@@
A.3 Prompts for Query-Expansion
@@
A.4 Details on Constructing Perspective Detection Test Set
@@
A.5 Details of Fine-tuning Mistral-7B Models
We fine-tune the Mistral-7B model with LoRA Hu et al. (2021) on eight NVIDIA A40 GPUs. We train the model on 3K examples of GPT-4 labeled data on perspective detection. The training dataset is constructed using the same process as described in Section A.4, except on 50 questions, yielding five times more data.
@@ A.5 Details of Fine-tuning Mistral-7B Models
Since the training examples are constructed using the same retriever-corpus settings, we also report results on out-of-domain examples in Section B.3.
A.6 Detail of Re-ranking Experiments
@@
A.7 Examples of Generated Perspectives for Query Expansion
@@
A.8 Details of Determining Supporting and Opposing Perspectives
@@
Tart 65.44 59.87 56.45 44.80 63.08 41.33 61.66 48.67
NV-Embed-v2 68.03 64.00 64.74 54.39 60.63 48.07 64.47 55.49
BM25 33.0 36.40 34.92 38.50 33.81 29.25 33.91 34.72
Google Search Contriever 48.65 44.53 47.95 44.32 42.73 35.49 46.44 41.45
Output Tart 51.43 45.33 49.99 43.20 45.69 35.07 49.04 41.20
NV-Embed-v2 59.17 55.87 61.61 51.60 50.32 40.13 57.03 49.20
@@
Appendix B Additional Results
B.1 Results for Top 10 Documents
@@
Corpus Model Avg. size ($\downarrow$) % within top 100 ($\uparrow$)
Wiki BM25 29.8 41.5
DPR 30.9 36.7
Cont. 24.7 55.6
Sphere BM25 20.8 77.3
DPR 37.4 48.9
Cont. 17.8 83.1
Google BM25 21.0 72.1
Cont. 18.2 76.2
B.2 How many documents should one retrieve to cover diverse perspectives?
In Table 3, we have reported the performances of the retrieval outputs consisting of five documents and found that the model outputs do not cover diverse perspectives. If we consider up to the top 100 documents, can we cover diverse perspectives? What is the average number of documents needed to be retrieved to cover all perspectives? Table 17 reports the size of the retrieved document set needed to cover all the perspectives with base retrievers. We find that the best setting (Contriver with Sphere) can find all the perspectives within the top 100 document set 83.1% of the time, and for those examples, the average number of documents required to cover all the perspectives is 17.8. Other settings show worse performance, mostly reflecting performance reported in Table 3.
B.3 Additional Validation Results on Perspective Detection
The training examples for fine-tuning the perspective detection model share the (corpus, retriever) settings for evaluation. We thus consider the results in Table 4 to be in-domain. We conduct further evaluation on out-of-domain examples. We sample 300 (Perspective, Document) pairs per (corpus, retriever) setting. We label each pair with GPT-4, and compare the trained Mistral perspective detection model against GPT-4 predictions. The results are in Table 18. The Mistral model is trained with GPT-4 labeled examples from the bolded settings (Sphere+BM25, Sphere+DPR). Our results show that it generalizes to out-of-domain settings.
Corpus Retriever Acc. F1 % Positive
(Labeled
by GPT-4)
Wiki BM25 92.3 68.1 12.7
Wiki DPR 92.1 70.4 14.2
Wiki Contriever 91.1 71.6 16.3
Sphere BM25 94.0 83.0 13.3
Sphere DPR 97.0 60.9 12.7
Sphere Contriever 89.3 75.8 14.2
Google Search BM25 91.5 76.7 19.2
Output
Google Search Contriever 89.0 76.3 24.3
Output
Table 18: Out-of-domain evaluation for the trained perspective detection model. The settings used to train the perspective detection model are bolded. The out-of-domain performances are roughly comparable to in-domain performances.
B.4 Manual Quality Inspection of BERDS
We manually examine 20 random examples in each subset of BERDS. The number of errors in each subset is presented in Table 19. We find that only one example in Arguana contains an unfaithful generated question, and the question itself is a valid contentious question (Table 20). In other words, the minor error does not affect the validity of the generated data. Only one example has a minor error (see Table 21) in Kialo, and it is because the provided positive perspective has a slight misalignment with the provided question. None of the examples in OpinionQA contain errors. The errors are minor and one can even argue the generated data is valid. Thus, we can infer that the generated data are of high quality.
Arguana Kialo OpinionQA
1 1 0
Table 19: We manually examine 20 random examples of each subset in BERDS, and show the number of errors. Only one example contains a minor error in each of Arguana and Kialo.
Question (Generated) Perspectives (Generated) Document 1 (Provided) Document 2 (Provided)
Is unrestricted access to birth control beneficial for society? [’Unrestricted access to birth control is beneficial for society.’, ’Unrestricted access to birth control is not beneficial for society.’] Any body of values that claims to respect the rights of the individual must recognise the right of a woman to choose Even the doctrines of the Church accepts that pregnancy is not, in and of itself, a virtue – there is no compulsion to maximise the number of pregnancies; there is simply a disagreement about how they should be avoided. The Church recommends that couples may minimise the chance without ever making it impossible through a chemical or physical barrier. In some parts of the world a pregnancy, even one that is not planned, is seen as a time for joy – a blessing for the family that will lead to a new and happy life bringing pleasure to both parents, their society and the child. That ideal is very far from the experience of much of the world where a child is another mouth to feed on impossibly little income. For all too much of the world, that life will be cruel, nasty and short. In slums, favellas and barren wastes that life is likely to be one marked more by dysentery or diarrhea, malnutrition and misery than by the sanitised, idealised image promoted in the West. […] It is difficult to see how the life of anyone is improved by reducing sex to a cheap form of entertainment. Certainly not the unborn children and not the objectified women. Proposition is more than happy for women to take control of their own fertility – indeed we would go further and suggest that their boyfriends and husbands should do so as well. Recreational sex, within wedlock and during times of infertility removes all of these problems; a little planning and restraint achieves that aim. It also means that both parents need to show that they are responsible for the results; Op seems happy to say that people are uncontrollable beasts with no control over their desires – hardly an edifying concept.
Table 20: Error Case of Arguana. The generated question may not have grasped the nuance of the discussion of the two provided documents, but it is relevant to the subject (birth control) and is a valid contentious question.
Question (Provided) Positive Perspective (Provided) Negative Perspective (Generated)
Should There be a Universal Basic Income (UBI)? Wealthy countries should provide citizens with a universal basic income (UBI). Wealthy countries should not provide citizens with a universal basic income (UBI).
Table 21: The error case of Kialo. The generated negative perspective is not faithful to the provided question only because the provided positive perspective itself mentions “wealthy countries".